OpenStudy (18jonea):

Which formula does NOT describe the sequence 8, 6, 3, –1, –6, … ? an = –0.5n2 – 0.5n + 9 an = an – 1 – n where a1 = 8 an = –0.5n2 – n + 9 None of the formulas describe the sequence. @jabez177

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

hint: let's consider the third formula: \[\Large {a_n} = - \frac{{{n^2}}}{2} - n + 9\] Now, if I replace \(n=1\), I get: \[\Large {a_1} = - \frac{1}{2} - 1 + 9 = - \frac{3}{2} + 9 = \frac{{15}}{2}\] which is different from \(8\). So, what can you conclude?

OpenStudy (18jonea):

Im kinda confused about this @Michele_Laino

OpenStudy (michele_laino):

hint: the third formula can not describe the sequence, since its first term is not equal to the first term of the given sequence

OpenStudy (18jonea):

Oh so c is the answer?

OpenStudy (michele_laino):

yes!

OpenStudy (18jonea):

Thank You